OpenStudy (anonymous):

Essay; show all work. A community 5K run will award $50 to the winner. 45 people enter the race, and they each pay an entry fee of $20. Assuming they are all equally likely to win, what is a fair price for the competition? Round to the nearest cent.

OpenStudy (anonymous):

What do you mean by a fair price?

OpenStudy (anonymous):

what is a good/reasonble price

OpenStudy (anonymous):

For what?

OpenStudy (anonymous):

That's what I don't know. Just tell your teacher it is worded badly.

OpenStudy (anonymous):

There your go, good job for finding the answer. I believe fair price means how much each competitor SHOULD pay, as opposed to how much they DID pay.

OpenStudy (anonymous):

$50X(1/45)-$20X44/45=-18.44+20.00=$1.56 and $1.56 is fair

OpenStudy (anonymous):

Exactly, very good maths!

OpenStudy (anonymous):

thank you